OpenStudy (anonymous):

slope=3; (1,5) Write in standard form the equation of each line.

OpenStudy (saifoo.khan):

y = mx + b 5 = 3(1) + b b = 5 - 3 b = 2 y = mx + b y = 3x +2

OpenStudy (anonymous):

y = mx + b substituting, \[ y-5=3(x-1) \Rightarrow y = 3x +2 \]

OpenStudy (anonymous):

I need it in Ax+By=C equation

OpenStudy (anonymous):

impossible

OpenStudy (saifoo.khan):

-3x + y = 2

OpenStudy (anonymous):

\[3x-y=-2\]

OpenStudy (anonymous):

the answer is -3x+y = 2.